Convert 4.77 picojoules to electron volts.


1
Expert's answer
2022-02-06T14:34:39-0500

Given:

"4.77\:\rm pJ=4.77*10^{-12}\: J"

"1\: \rm eV=1.6*10^{-19}\: J"




"{4.77\:\rm pJ=4.77*10^{-12}\: J/(1.6*10^{-19}\: J/eV)}\\
=2.98*10^7\: \rm eV=29.8\: MeV"